# The leading constant in Malle’s conjecture over function fields

### v1

@@ 4.2.3. Volume computations
For all $c\in\pi_{0}(\mathcal{C}^{*}_{G})$ we introduce a formal variable $T_{c}$ and use the multi-index notation $\underline{T}^{\underline{n}}=\sum_{c}T_{c}^{n_{c}}$. For $\varphi_{P}\in BG[K_{P}]$, we write $T(\varphi_{P}):=T_{c}^{\deg P}$ such that $c$ is the ramification type of $\varphi_{P}$. For $(\varphi_{P})_{P}\in BG[\mathbf{A}_{K,\infty}]$, we write that $T(\varphi)=\prod_{P}T(\varphi_{P})$.
@@
Let $\beta\in\operatorname{Br}BG_{\mathbb{F}_{q}}$ and let $\mathcal{C}_{\beta}\subset\mathcal{C}_{G}^{*}$ be the set of conjugacy classes where $\beta$ has an algebraic residue. For $c\in\pi_{0}(\mathcal{C}_{\beta})$ with field of definition $\mathbb{F}_{q}(c)$ we thus have a residue $\partial_{c}(\beta)\in\operatorname{H}^{1}(\mathbb{F}_{q}(c),\mathbb{Q}/\mathbb{Z})$. We then have the equality
@@
Let $\beta\in\operatorname{Br}BG_{\mathbb{F}_{q}}$ and let $\mathcal{C}_{\beta}\subset\mathcal{C}_{G}^{*}$ be the set of conjugacy classes where $\beta$ has an algebraic residue.
@@
    $\hat{\tau}(\beta;BG[\mathbf{A}_{\mathbb{F}_{q}(t)}]_{\underline{n},\Omega_{\infty}})=\hat{\tau}_{\infty}(\beta;\Omega_{\infty})\#\mathrm{Conf}_{\underline{n}}(\mathbb{F}_{q})\prod_{c\in\pi_{0}(\mathcal{C}_{\beta})}e(n_{c}\mathrm{cor}_{\mathbb{F}_{q}(c)/\mathbb{F}_{q}}(\partial_{c}(\beta))).$
@@
For all $c\in\pi_{0}(\mathcal{C}_{G})$ we introduce a formal variable $T_{c}$ and use the multi-index notation $\underline{T}^{\underline{n}}=\sum_{c}T_{c}^{n_{c}}$.
@@
$\hat{\tau}(\beta;\underline{n})\underline{T}^{\underline{n}}=\#\mathrm{Conf}_{\underline{n}}(\mathbb{F}_{q})\hat{\underline{T}}^{\underline{n}}=\#\mathrm{Conf}_{\underline{n}}(\mathbb{F}_{q})\prod_{c\in\pi_{0}(\mathcal{C}_{\beta})}e(\partial_{c}(\beta))\underline{T}^{\underline{n}}$
@@ 4.3. Multi-height counting
$\tau_{BG,\mathbb{F}_{q}}(BG[\mathbf{A}_{\mathbb{F}_{q}(t)}]_{\underline{n},\Omega_{\infty}}^{\operatorname{Br}})\leq\tau_{BG,\mathbb{F}_{q}}(BG[\mathbf{A}_{\mathbb{F}_{q}(t)}]_{\underline{n},\Omega_{\infty}})=\tau_{\infty}(\Omega_{\infty})\mathrm{Conf}_{\underline{n}}(\mathbb{F}_{q})\leq\tau_{\infty}(\Omega_{\infty})q^{|n|}$
@@
$N(G,\underline{n},\Omega_{\infty})=\frac{|Z(G(\mathbb{F}_{q}))|}{|G^{\mathrm{ab}}(-1)(\mathbb{F}_{q})|}\sum_{\beta\in\operatorname{Br}_{\bar{\mathcal{C}}}BG_{\bar{\mathbb{F}}_{q}}}\hat{\tau}(\beta;BG[\mathbf{A}_{\mathbb{F}_{q}(t)}]_{\underline{n},\Omega_{\infty}}).$
@@ 4.4.3. Unbalanced height
Note that this also holds if the support of $\underline{n}$ does not contain $\mathcal{C}_{f}$ as the error term dominates the main term by the same argument as in the proof of Theorem˜4.10.
The lemma then follows by summing of $\underline{n}$ and switching the order of summation. The sum over error terms is dealt with via Lemma˜4.12. ∎
It thus remains to deal for all $M\subseteq L\subseteq G$ with the sum

### v2

@@ 4.2.3. Volume computations
For all $c\in\pi_{0}(\mathcal{C}^{*}_{G})$ we introduce a formal variable $T_{c}$ and use the multi-index notation $\underline{T}^{\underline{n}}=\prod_{c}T_{c}^{n_{c}}$. For $\varphi_{P}\in BG[K_{P}]$, we write $T(\varphi_{P}):=T_{c}^{\deg P}$ such that $c$ is the ramification type of $\varphi_{P}$. For $(\varphi_{P})_{P}\in BG[\mathbf{A}_{K,\infty}]$, we write that $T(\varphi)=\prod_{P}T(\varphi_{P})$.
@@
Let $\beta\in\operatorname{Br}BG_{\mathbb{F}_{q}}$ and let $\mathcal{C}_{\beta}\subset\mathcal{C}_{G}^{*}$ be the set of conjugacy classes $c$ where the residue of $\beta$ at $c$ is algebraic. For $c\in\pi_{0}(\mathcal{C}_{\beta})$ with field of definition $\mathbb{F}_{q}(c)$ we thus have $\partial_{c}(\beta)\in\operatorname{H}^{1}(\mathbb{F}_{q}(c),\mathbb{Q}/\mathbb{Z})$. We then have the equality
@@
Let $\beta\in\operatorname{Br}BG_{\mathbb{F}_{q}}$ and let $\mathcal{C}_{\beta}\subset\mathcal{C}_{G}^{*}$ be the set of conjugacy classes $c$ where the residue $\partial_{c}(\beta)$ is algebraic.
@@
    $\hat{\tau}(\beta;BG[\mathbf{A}_{\mathbb{F}_{q}(t)}]_{\underline{n},\Omega_{\infty}})=\hat{\tau}_{\infty}(\beta;\Omega_{\infty})\#\mathrm{Conf}_{\mathcal{C}_{\beta},\underline{n}}(\mathbb{F}_{q})\prod_{c\in\pi_{0}(\mathcal{C}_{\beta})}e(n_{c}\mathrm{cor}_{\mathbb{F}_{q}(c)/\mathbb{F}_{q}}(\partial_{c}(\beta))).$
@@
For all $c\in\pi_{0}(\mathcal{C}_{G})$ we introduce a formal variable $T_{c}$ and use the multi-index notation $\underline{T}^{\underline{n}}=\prod_{c}T_{c}^{n_{c}}$.
@@
$\hat{\tau}(\beta;\underline{n})\underline{T}^{\underline{n}}=\#\mathrm{Conf}_{\mathcal{C}_{\beta},\underline{n}}(\mathbb{F}_{q})\hat{\underline{T}}^{\underline{n}}=\#\mathrm{Conf}_{\mathcal{C}_{\beta},\underline{n}}(\mathbb{F}_{q})\prod_{c\in\pi_{0}(\mathcal{C}_{\beta})}e(\partial_{c}(\beta))\underline{T}^{\underline{n}}$
@@ 4.3. Multi-height counting
$\tau_{BG,\mathbb{F}_{q}}(BG[\mathbf{A}_{\mathbb{F}_{q}(t)}]_{\underline{n},\Omega_{\infty}}^{\operatorname{Br}})\leq\tau_{BG,\mathbb{F}_{q}}(BG[\mathbf{A}_{\mathbb{F}_{q}(t)}]_{\underline{n},\Omega_{\infty}})=\tau_{\infty}(\Omega_{\infty})\mathrm{Conf}_{\mathcal{C}_{G}^{*},\underline{n}}(\mathbb{F}_{q})\leq\tau_{\infty}(\Omega_{\infty})q^{|n|}$
@@
$N(G,\underline{n},\Omega_{\infty})=\frac{|Z(G(\mathbb{F}_{q}))|}{|G^{\mathrm{ab}}(-1)(\mathbb{F}_{q})|}\sum_{\beta\in\operatorname{Br}_{\bar{\mathcal{C}}}BG_{\bar{\mathbb{F}}_{q}}}\hat{\tau}(\beta;BG[\mathbf{A}_{\mathbb{F}_{q}(t)}]_{\underline{n},\Omega_{\infty}})+O(q^{|\underline{n}|-\delta\min_{c\in\mathcal{C}_{\mathrm{st}}(\bar{\mathbb{F}}_{q})}(n_{c})}).$
@@ 4.4.3. Unbalanced heights
Note that this also holds if the support of $\underline{n}$ does not contain $\mathcal{C}_{f}$ as in this case the error term dominates the main term by the same argument as in the proof of Theorem˜4.10.
The lemma then follows by summing over $\underline{n}$ and switching the order of summation. The sum over error terms is dealt with in Lemma˜4.12. ∎
It thus remains to estimate for all $M\subseteq L\subseteq G$ the sum